# Hao Shi1  Bin Xie2  Yingfei Liu2  Lin Sun4  Fengrong Liu5  Tiancai Wang2 Erjin Zhou2  Haoqiang Fan2  Xiangyu Zhang3,6  Gao Huang1🖂

### v1

@@ Abstract
Temporal context is essential for robotic manipulation because such tasks are inherently non-Markovian, yet mainstream VLA models typically overlook it and struggle with long-horizon, temporally dependent tasks. Cognitive science suggests that humans rely on working memory to buffer short-lived representations for immediate control, while the hippocampal system preserves verbatim episodic details and semantic gist of past experience for long-term memory. Inspired by these mechanisms, we propose MemoryVLA, a Cognition-Memory-Action framework for long-horizon robotic manipulation. A pretrained VLM encodes the observation into perceptual and cognitive tokens that form working memory, while a Perceptual-Cognitive Memory Bank stores low-level details and high-level semantics consolidated from it. Working memory retrieves decision-relevant entries from the bank, adaptively fuses them with current tokens, and updates the bank by merging redundancies. Using these tokens, a memory-conditioned diffusion action expert yields temporally aware action sequences. We evaluate MemoryVLA on 150+ simulation and real-world tasks across three robots. On SimplerEnv-Bridge, Fractal, and LIBERO-5 suites, it achieves 71.9%, 72.7%, and 96.5% success rates, respectively, all outperforming state-of-the-art baselines CogACT and $\pi_{0}$, with a notable +14.6 gain on Bridge. On 12 real-world tasks spanning general skills and long-horizon temporal dependencies, MemoryVLA achieves 84.0% success rate, with long-horizon tasks showing a +26 improvement over state-of-the-art baseline.  Project Page is available.
@@ 1 Introduction
Vision-Language-Action (VLA) models (Brohan et al., 2023; Kim et al., 2024; Black et al., 2024; Li et al., 2024a; Kim et al., 2025; Sun et al., 2025), powered by large-scale cross-embodiment robotic datasets (O’Neill et al., 2024; Walke et al., 2023; Brohan et al., 2022; Khazatsky et al., 2024; Bu et al., 2025a) and pretrained Vision-Language Models (VLMs) (Karamcheti et al., 2024; Liu et al., 2023b; Bai et al., 2023a), have achieved remarkable progress in robotic manipulation. However, mainstream VLA models such as OpenVLA (Kim et al., 2024) and $\pi_{0}$ (Black et al., 2024) rely solely on the current observation, thereby overlooking temporal dependencies and performing poorly on long-horizon temporal manipulation tasks. As shown in Fig. 1 (a), Push Buttons tasks exhibit almost no visual difference before and after pushing, making it difficult to determine whether the action has already been completed. This highlights the non-Markovian nature of manipulation, where earlier actions influence later decisions, calling for temporal modeling. A naive strategy is to concatenate consecutive frames as input to the VLM. However, it faces two critical limitations: (1) The quadratic complexity of self-attention severely limits the usable temporal context length; (2) Sequential frame inputs are misaligned with the model’s single-frame robotic pretraining distribution.
@@
Research in cognitive science (Baddeley & Hitch, 1974; Tulving et al., 1972; Reyna & Brainerd, 1995) demonstrates that humans handle manipulation tasks through a dual-memory system (Fig. 1 (b)). The brain encodes multi-modal sensory inputs into both perceptual and cognitive representations. These representations are buffered in working memory via transient neural activity, providing short-term retention for immediate decision-making. Concurrently, episodic memory, the long-term memory system supported by hippocampus, encodes past experiences with temporal index in two forms: verbatim representations preserving precise details and gist representations capturing abstract semantics. During execution, working memory retrieves decision-relevant contexts from episodic memory and integrates them with current representations to guide actions through cerebellar control, while simultaneously consolidating new experiences into episodic memory.
@@
We conduct extensive evaluations of MemoryVLA across 3 robots and 150+ tasks with 500+ variations in simulation and real world. For SimplerEnv (Li et al., 2024b), MemoryVLA achieves 71.9% and 72.7% success rates on Bridge and Fractal suites, surpassing CogACT by 14.6 and 4.6 points, further outperforming $\pi_{0}$. For LIBERO (Liu et al., 2023a), it achieves an average success rate of 96.5% across 5 suites (Spatial, Object, Goal, Long, and LIBERO-90), exceeding both CogACT and $\pi_{0}$. For real-world evaluations, we introduce 12 tasks across Franka and WidowX robots, spanning 6 general tasks and 6 long-horizon temporal tasks. MemoryVLA achieves 85% and 83% scores on general and temporal tasks, outperforming CogACT by 9 and 26 points, and substantially surpassing $\pi_{0}$. Moreover, MemoryVLA exhibits strong robustness and generalization under out-of-distribution conditions involving varied backgrounds, distractors, objects, containers, lighting and occlusion.
@@
  • MemoryVLA achieves state-of-the-art performance on SimplerEnv, LIBERO, and real-world. It also demonstrates strong robustness and generalization. On challenging long-horizon real-world tasks, it outperforms CogACT and $\pi_{0}$ by significant margins, underscoring the importance of temporal memory modeling.
@@ 2 Related Works
Driven by advances in visual foundation models (Radford et al., 2021; Caron et al., 2021; Liu et al., 2024a; Zheng et al., 2024a; 2025a; Zhang et al., 2025b), robot imitation learning has progressed rapidly yet remains confined to small, task-specific policies with limited generalization (Shridhar et al., 2023; Zhao et al., 2023; Chi et al., 2023; Goyal et al., 2023). To overcome these, the success of VLMs (Achiam et al., 2023; Touvron et al., 2023; Liu et al., 2023b; Bai et al., 2023b) and large-scale robot datasets (e.g., OXE (O’Neill et al., 2024), Agibot (Bu et al., 2025a)) spawned the vision-language-action (VLA) paradigm. RT-2 (Zitkovich et al., 2023) and OpenVLA (Kim et al., 2024) tokenize continuous actions into discrete tokens and use VLMs for autoregressive prediction as if generating language. In contrast, $\pi_{0}$ (Black et al., 2024), CogACT (Li et al., 2024a), DexVLA (Wen et al., 2025) and HybridVLA (Liu et al., 2025b) adopt diffusion-based policies (Chi et al., 2023; Liu et al., 2024b) as action heads, leveraging iterative denoising to sample continuous control trajectories that capture diverse multimodal behaviors. However, none of these methods explicitly model temporal dependencies. Robotic manipulation is inherently non-Markovian, and neglecting history leads to failures on long-horizon temporal tasks.
@@
Temporal modeling has been extensively studied in computer vision and autonomous driving (Wang et al., 2023; Liu et al., 2023c; Feng et al., 2023; Zhou et al., 2024), yet it has not been fully explored in robotic manipulation. Octo (Mees et al., 2024), RoboVLMs (Liu et al., 2025a), and Interleave-VLA (Fan et al., 2025) adapt the VLM paradigm to model robotic video data in an interleaved image-text format. While conceptually elegant, this format is complex to implement and computationally expensive, hindering its widespread application. RoboFlamingo (Li et al., 2023) compresses vision-language representation into a latent token and propagate it via LSTM (Hochreiter & Schmidhuber, 1997). The latent representation is obtained in a relatively coarse manner and the fine-grained perceptual history is largely discarded. TraceVLA (Zheng et al., 2024b) takes a different route, painting historical states as trajectories on the current frame, yet discards rich semantic details. UniVLA (Bu et al., 2025b) incorporates past actions into input prompts, making an initial attempt at temporal modeling. However, it merely serves as a Chain-of-Thought (Wei et al., 2022) process without effectively utilizing historical information. In contrast, we model both high-level cognitive semantics and fine-grained perceptual details within a memory framework, enabling effective temporal modeling for long-horizon manipulation.
@@ 3.2 Vision-Language Cognition Module
We build upon a 7B–parameter Prismatic VLM (Karamcheti et al., 2024), which is further pretrained on the large-scale cross-embodiment real robot dataset Open-X Embodiment (O’Neill et al., 2024). For visual encoding, we adopt parallel DINOv2 (Oquab et al., 2023) and SigLIP (Zhai et al., 2023) backbones on the current third-person RGB image $I$, concatenating their features into raw visual tokens. A perceptual compression module, implemented via a SE-bottleneck (Hu et al., 2018), then compresses these tokens into a compact set of perceptual tokens $p\in\mathbb{R}^{N_{p}\times d_{p}}$ with $N_{p}=256$. In parallel, the raw visual tokens are projected via a linear layer into the language embedding space and concatenated with the tokenized instruction before being fed into the LLaMA-7B (Touvron et al., 2023). The output at the end-of-sentence (EOS) position is taken as the cognitive token $c\in\mathbb{R}^{1\times d_{c}}$, representing high-level cognitive semantics in compact form. Finally, the perceptual tokens $p$ and cognitive token $c$ are combined to form the short-term working memory for downstream modules.
@@ 3.3 Perceptual-Cognitive Memory Module
This consolidation mechanism mitigates memory bloat by reducing redundancy, while preserving the most salient perceptual details and semantic abstractions, thereby maintaining a compact representation that supports efficient long-term memory.
@@ 3.4 Memory-Conditioned Action Expert
Leveraging the memory-augmented working memory $\{\tilde{p},\tilde{c}\}$, which integrates historical perceptual and cognitive information, the action expert predicts a sequence of future actions $\{a_{1},a_{2},\dots,a_{T}\}$, with $T=16$. This prediction allows the model to anticipate multi-step trajectories, reduce cumulative error, and provide foresight for long-horizon execution. Since real-world robotic actions lie in a continuous multimodal control space, we adopt a diffusion-based Transformer (DiT) (Peebles & Xie, 2023) implemented with Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020), using 10 denoising steps for efficient yet accurate trajectory generation. This architecture progressively denoises a sequence of noisy action tokens, yielding precise continuous actions.
@@
Figure 4: Experimental setup overview. Top: three simulation benchmarks, SimpleEnv-Bridge with WidowX, SIMPLER-Fractal with Google Robot, and LIBERO with Franka. Bottom: real-world evaluation on two suites, General and Long-horizon Temporal. In total, we evaluate three robots across 10 suites, spanning over 150 tasks and 500 variations.
@@ 4 Experiments
To comprehensively evaluate MemoryVLA, we organize experiments around five core questions: (1) How does MemoryVLA compare with state-of-the-art methods on SimplerEnv? (Sec. 4.2) (2) How does it perform on LIBERO? (Sec. 4.3) (3) Can it handle both general manipulation and long-horizon temporal tasks on real robots? (Sec. 4.4) (4) What is the impact of each component? (Sec. 4.5) (5) How robust and generalizable is it under diverse environmental conditions? (Appendix A)
Table 1: Performance comparison on SimplerEnv-Bridge (Li et al., 2024b) with WidowX robot. CogACT-Large is our re-evaluated baseline using official weight, and MemoryVLA achieves a +14.6 gain in average success. Entries marked with * are reproduced from open-pi-zero, which leverage additional proprioceptive state inputs; they also adopt Uniform/Beta timestep sampling.
Method
Spoon
on Towel
Carrot
on Plate
Stack
Cube
Eggplant
in Basket
Avg.
Success
RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1
OpenVLA (Kim et al., 2024) 4.2 0.0 0.0 12.5 4.2
Octo-Base (Team et al., 2024) 15.8 12.5 0.0 41.7 17.5
TraceVLA (Zheng et al., 2024b) 12.5 16.6 16.6 65.0 27.7
RoboVLMs (Liu et al., 2025a) 45.8 20.8 4.2 79.2 37.5
SpatialVLA (Qu et al., 2025) 16.7 25.0 29.2 100.0 42.7
Magma (Yang et al., 2025) 37.5 29.2 20.8 91.7 44.8
CogACT-Base (Li et al., 2024a) 71.7 50.8 15.0 67.5 51.3
$\pi_{0}$-Uniform* (Black et al., 2024) 63.3 58.8 21.3 79.2 55.7
CogACT-Large (Li et al., 2024a) 58.3 45.8 29.2 95.8 57.3
$\pi_{0}$-Beta* (Black et al., 2024) 84.6 55.8 47.9 85.4 68.4
MemoryVLA (Ours) 75.0 75.0 37.5 100.0 71.9(+14.6)
Table 2: Performance comparison on SimplerEnv-Fractal (Li et al., 2024b) with Google robot. Success rates (%) are reported for Visual Matching (VM) and Visual Aggregation (VA) suites. MemoryVLA achieves an overall +4.6 gain over CogACT. O./C. denotes Open/Close, and * follow Tab. 1.
Method
Visual Matching (VM) Visual Aggregation (VA)
Overall
Coke
Can
Move
Near
O. / C.
Drawer
Put in
Drawer
Avg.
Coke
Can
Move
Near
O. / C.
Drawer
Put in
Drawer
Avg.
Octo-Base (Team et al., 2024) 17.0 4.2 22.7 0.0 11.0 0.6 3.1 1.1 0.0 1.2 6.1
RT-1-X (O’Neill et al., 2024) 56.7 31.7 59.7 21.3 42.4 49.0 32.3 29.4 10.1 30.2 36.3
OpenVLA (Kim et al., 2024) 18.0 56.3 63.0 0.0 34.3 60.8 67.7 28.8 0.0 39.3 36.8
RoboVLMs (Liu et al., 2025a) 76.3 79.0 44.9 27.8 57.0 50.7 62.5 10.3 0.0 30.9 44.0
TraceVLA (Zheng et al., 2024b) 45.0 63.8 63.1 11.1 45.8 64.3 60.6 61.6 12.5 49.8 47.8
RT-2-X (O’Neill et al., 2024) 78.7 77.9 25.0 3.7 46.3 82.3 79.2 35.5 20.6 54.4 50.4
Magma (Yang et al., 2025) 75.0 53.0 58.9 8.3 48.8 68.6 78.5 59.0 24.0 57.5 53.2
SpatialVLA (Qu et al., 2025) 79.3 90.0 54.6 0.0 56.0 78.7 83.0 39.2 6.3 51.8 53.9
$\pi_{0}$-Uniform* (Black et al., 2024) 88.0 80.3 56.0 52.2 69.1
$\pi_{0}$-Beta* (Black et al., 2024) 97.9 78.7 62.3 46.6 71.4
CogACT (Li et al., 2024a) 91.3 85.0 71.8 50.9 74.8 89.6 80.8 28.3 46.6 61.3 68.1
@@ 4.1 Experimental Setups
Fig. 4 overviews our evaluation across simulation and real-world, covering 3 robots, 10 suites, 150+ tasks with 500+ variations. SimplerEnv (Li et al., 2024b) includes Bridge suite with a WidowX robot and Fractal suite with a Google robot. Fractal provides two settings: Visual Matching (VM) and Visual Aggregation (VA). LIBERO (Liu et al., 2023a) uses a Franka robot and spans five suites (Spatial, Object, Goal, Long, and LIBERO-90). In real-world, we evaluate General and Long-horizon Temporal suites on Franka and WidowX robots.
@@
We train on 8 NVIDIA A100 GPUs with PyTorch FSDP, using 32 samples per GPU for a global batch of 256 and a learning rate of $2\times 10^{-5}$. The model takes a single third-person RGB frame at $224\times 224$ together with the language instruction and outputs 7-DoF actions. The LLM is 7B, and the diffusion action expert has $\sim$300M parameters. At inference we use DDIM (Song et al., 2020) with 10 sampling steps and a classifier-free guidance(CFG) (Ho & Salimans, 2022) guidance scale of 1.5. Additional details are provided in Appendix B and C.

### v2

@@ Abstract
Temporal context is essential for robotic manipulation because such tasks are inherently non-Markovian, yet mainstream VLA models typically overlook it and struggle with long-horizon, temporally dependent tasks. Cognitive science suggests that humans rely on working memory to buffer short-lived representations for immediate control, while the hippocampal system preserves verbatim episodic details and semantic gist of past experience for long-term memory. Inspired by these mechanisms, we propose MemoryVLA, a Cognition-Memory-Action framework for long-horizon robotic manipulation. A pretrained VLM encodes the observation into perceptual and cognitive tokens that form working memory, while a Perceptual-Cognitive Memory Bank stores low-level details and high-level semantics consolidated from it. Working memory retrieves decision-relevant entries from the bank, adaptively fuses them with current tokens, and updates the bank by merging redundancies. Using these tokens, a memory-conditioned diffusion action expert yields temporally aware action sequences. We evaluate MemoryVLA on 150+ simulation and real-world tasks across three robots. On SimplerEnv-Bridge, Fractal, LIBERO-5 suites and Mikasa-Robo, it achieves 71.9%, 72.7%, 96.5%, and 41.2% success rates, respectively, all outperforming state-of-the-art baselines CogACT and $\pi_{0}$, with a notable +14.6 gain on Bridge and +11.8 gain on Mikasa-Robo. On 12 real-world tasks spanning general skills and long-horizon temporal dependencies, MemoryVLA achieves 84.0% success score, with long-horizon tasks showing a +26 improvement over state-of-the-art baseline. Project Page: https://shihao1895.github.io/MemoryVLA
@@ 1 Introduction
Vision-Language-Action (VLA) models (brohan2023rt; kim2024openvla; black2024pi_0; li2024cogact; sun2025geovla; xie2025dexbotic), powered by large-scale cross-embodiment robotic datasets (o2024open; brohan2022rt; khazatsky2024droid; bu2025agibot) and pretrained Vision-Language Models (VLMs) (karamcheti2024prismatic; liu2023visual; qwen), have achieved remarkable progress in robotic manipulation. However, mainstream VLA models such as OpenVLA (kim2024openvla) and $\pi_{0}$ (black2024pi_0) rely solely on the current observation, thereby overlooking temporal dependencies and performing poorly on long-horizon temporal manipulation tasks. As shown in Fig. 1 (a), Push Buttons tasks exhibit almost no visual difference before and after pushing, making it difficult to determine whether the action has already been completed. This highlights the non-Markovian nature of manipulation, where earlier actions influence later decisions, calling for temporal modeling. A naive strategy is to concatenate consecutive frames as input to the VLM. However, it faces two critical limitations: (1) The quadratic complexity of self-attention severely limits the usable temporal context length; (2) Sequential frame inputs are misaligned with the model’s single-frame robotic pretraining distribution.
@@
Research in cognitive science (baddeley1974working; tulving1972episodic; reyna1995fuzzy) demonstrates that humans handle manipulation tasks through a dual-memory system (Fig. 1 (b)). The brain encodes multi-modal sensory inputs into both perceptual and cognitive representations. These representations are buffered in working memory via transient neural activity, providing short-term retention for immediate decision-making. Concurrently, episodic memory, the long-term memory system supported by hippocampus, encodes past experiences with temporal index in two forms: verbatim representations preserving precise details and gist representations capturing abstract semantics. During execution, working memory retrieves decision-relevant contexts from episodic memory and integrates them with current representations to guide actions through cerebellar control, while simultaneously consolidating new experiences into episodic memory.
@@
We conduct extensive evaluations of MemoryVLA across 3 robots and 150+ tasks with 500+ variations in simulation and real world. For SimplerEnv (li2024evaluating), MemoryVLA achieves 71.9% and 72.7% success rates on Bridge and Fractal suites, surpassing CogACT by 14.6 and 4.6 points, further outperforming $\pi_{0}$. For LIBERO (liu2023libero), it achieves 96.5% success rate across 5 suites (Spatial, Object, Goal, Long-10, and Long-90), exceeding both CogACT and $\pi_{0}$. For Mikasa-Robo (cherepanov2025memory), it achieves 41.2% success rate, outperforming $\pi_{0}$ by 11.8 points. For real-world evaluations, we introduce 12 tasks across Franka and WidowX robots, spanning 6 general tasks and 6 long-horizon temporal tasks. MemoryVLA achieves 85% and 83% scores on general and temporal tasks, outperforming CogACT by 9 and 26 points, and substantially surpassing $\pi_{0}$. Moreover, MemoryVLA exhibits strong robustness and generalization under out-of-distribution conditions involving varied backgrounds, distractors, objects, containers, lighting and occlusion.
@@
  • MemoryVLA achieves state-of-the-art performance on SimplerEnv, LIBERO, Mikasa-Robo, and real-world. It also demonstrates strong robustness and generalization. On challenging long-horizon real-world tasks, it outperforms CogACT and $\pi_{0}$ by significant margins, underscoring the importance of temporal memory modeling.
@@ 2 Related Works
Driven by advances in visual foundation models (radford2021learning; caron2021emerging; liu2024grounding; zheng2024denseg; zheng2025densegrounding; zhang2025grounding; wu2025ragnet; wang2025emulating), robot imitation learning has progressed rapidly yet remains confined to small, task-specific policies with limited generalization (shridhar2023perceiver; zhao2023learning; chi2023diffusion; goyal2023rvt; shi2025spatialactor). To overcome these, the success of VLMs (achiam2023gpt; touvron2023llama; liu2023visual; bai2023qwen; guo2025deepseek) and large-scale robot datasets (e.g., OXE (o2024open), Agibot (bu2025agibot)) spawned the vision-language-action (VLA) paradigm (kim2024openvla; black2024pi_0; yue2024deer; zhong2025survey; gao2025vla; yang2025instructvla). RT-2 (zitkovich2023rt) and OpenVLA (kim2024openvla) tokenize continuous actions into discrete tokens and use VLMs for autoregressive prediction as if generating language. In contrast, $\pi_{0}$ (black2024pi_0), CogACT (li2024cogact), DexVLA (wen2025dexvla) and HybridVLA (liu2025hybridvla) adopt diffusion-based policies (chi2023diffusion; liu2024rdt) as action heads, leveraging iterative denoising to sample continuous control trajectories that capture diverse multimodal behaviors. However, none of these methods explicitly model temporal dependencies. Robotic manipulation is inherently non-Markovian, and neglecting history leads to failures on long-horizon temporal tasks.
@@
Temporal modeling has been extensively studied in computer vision and autonomous driving (wang2023exploring; liu2023petrv2; feng2023open; zhou2024rmem), yet it has not been fully explored in robotic manipulation. Octo (mees2024octo), RoboVLMs (liu2025towards), and Interleave-VLA (fan2025interleave) adapt the VLM paradigm to model robotic video data in an interleaved image-text format. While conceptually elegant, this format is complex to implement and computationally expensive, hindering its widespread application. RoboFlamingo (li2023vision) compresses vision-language representation into a latent token and propagate it via LSTM (hochreiter1997long). The latent representation is obtained in a relatively coarse manner and the fine-grained perceptual history is largely discarded. TraceVLA (zheng2024tracevla) takes a different route, painting historical states as trajectories on the current frame, yet discards rich semantic details. UniVLA (bu2025univla) incorporates past actions into input prompts, making an initial attempt at temporal modeling. However, it merely serves as a Chain-of-Thought (wei2022chain) process without effectively utilizing historical information. In contrast, we model both high-level cognitive semantics and fine-grained perceptual details within a memory framework, enabling effective temporal modeling for long-horizon manipulation.
@@ 3.2 Vision-Language Cognition Module
We build upon a 7B–parameter Prismatic VLM (karamcheti2024prismatic), which is further pretrained on the large-scale cross-embodiment real robot dataset Open-X Embodiment (o2024open). For visual encoding, we adopt parallel DINOv2 (oquab2023dinov2) and SigLIP (zhai2023sigmoid) backbones on the current third-person RGB image $I$, concatenating their features into raw visual tokens. A perceptual compression module, implemented via a SE-bottleneck (hu2018squeeze), then compresses these tokens into a compact set of perceptual tokens $p\in\mathbb{R}^{N_{p}\times d_{p}}$ with $N_{p}=256$. In parallel, the raw visual tokens are projected via a linear layer into the language embedding space and concatenated with the tokenized instruction before being fed into the LLaMA-7B (touvron2023llama). The output at the end-of-sentence (EOS) position is taken as the cognitive token $c\in\mathbb{R}^{1\times d_{c}}$, representing high-level cognitive semantics in compact form. Finally, the perceptual tokens $p$ and cognitive token $c$ are combined to form the short-term working memory for downstream modules.
@@ 3.3 Perceptual-Cognitive Memory Module
This consolidation mechanism (Fig. 3 (c)) mitigates memory bloat by reducing redundancy, while preserving the most salient perceptual details and semantic abstractions, thereby maintaining a compact representation that supports efficient long-term memory.
@@ 3.4 Memory-Conditioned Action Expert
Leveraging the memory-augmented working memory $\{\tilde{p},\tilde{c}\}$, which integrates historical perceptual and cognitive information, the action expert predicts a sequence of future actions $\{a_{1},a_{2},\dots,a_{T}\}$, with $T=16$. This prediction allows the model to anticipate multi-step trajectories, reduce cumulative error, and provide foresight for long-horizon execution. Since real-world robotic actions lie in a continuous multimodal control space, we adopt a diffusion-based Transformer (DiT) (peebles2023scalable) implemented with Denoising Diffusion Implicit Models (DDIM) (song2020denoising), using 10 denoising steps for efficient yet accurate trajectory generation. This architecture progressively denoises a sequence of noisy action tokens, yielding precise continuous actions.
@@ 4 Experiments
To comprehensively evaluate MemoryVLA, we organize experiments around six core questions: (1) How does MemoryVLA compare with state-of-the-art methods on SimplerEnv benchmark? (Sec. 4.2) (2) How does it perform on LIBERO benchmark? (Sec. 4.3) (3) How does it perform on Mikasa-Robo benchmark? (Sec. 4.4) (4) Can it handle both general manipulation and long-horizon temporal tasks on real robots? (Sec. 4.5) (5) What is the impact of each component? (Sec. 4.6) (6) How robust and generalizable is it under diverse environmental conditions? (Appendix B)
@@ 4.1 Experimental Setups
Fig. 4 overviews our evaluation across simulation and real-world, covering 3 robots, 6 benchmarks, 150+ tasks with 500+ variations. SimplerEnv (li2024evaluating) includes Bridge suite with a WidowX robot and Fractal suite with a Google robot. Fractal provides two settings: Visual Matching (VM) and Visual Aggregation (VA). LIBERO (liu2023libero) uses a Franka robot and spans five suites (Spatial, Object, Goal, Long-10, and Long-90). Mikasa-Robo (cherepanov2025memory) uses a Franka robot. In real-world, we evaluate General and Long-horizon Temporal suites on Franka and WidowX robots. Task details for each benchmark and additional qualitative results are provided in Appendix K and Appendix M.
@@
We train on 8 NVIDIA A100 GPUs with PyTorch FSDP, using 32 samples per GPU for a global batch of 256 and a learning rate of $2\times 10^{-5}$. The model takes a single third-person RGB frame at $224\times 224$ together with the language instruction and outputs 7-DoF actions. The LLM is 7B, and the diffusion action expert has $\sim$300M parameters. At inference we use DDIM (song2020denoising) with 10 sampling steps and a classifier-free guidance(CFG) (ho2022classifier) guidance scale of 1.5. Additional details are provided in Appendix C and D.
Figure 4: Experimental setup overview. Top: Four simulation benchmarks, SimpleEnv-Bridge, SimpleEnv-Fractal, LIBERO, and Mikasa-Robo. Bottom: real-world evaluation (General and Long-horizon Temporal), real-world robustness and generalization evaluation. In total, we evaluate 3 robots across 6 benchmarks, spanning over 150 tasks and 500 variations.